OpenStudy (anonymous):

The surface area of a rectangular prism is 15 m2. What is the surface area of this prism, measured in cm2?

OpenStudy (ipwnbunnies):

Do you know how many centimeters are in a meter?

OpenStudy (anonymous):

10 or 20

OpenStudy (ipwnbunnies):

No, there are 100 centimeters in a meter.

OpenStudy (ipwnbunnies):

\[(15 m^{2})(\frac{100 cm}{1 m})^{2}\]

OpenStudy (ipwnbunnies):

100/1 is our conversion factor, because we need centimeters. So we eliminate the meters.

OpenStudy (ipwnbunnies):

BUT, we also need SQUARE centimeters for the area. So, we also square the conversion factor.

OpenStudy (anonymous):

A. 0.0015 cm2 B. 0.15 cm2 C. 150,000 cm2 D. 1500 cm2

OpenStudy (ipwnbunnies):

What'd you get when you did the math? Make sure to square the conversion factor first.

OpenStudy (anonymous):

a or b

OpenStudy (ipwnbunnies):

Neither...Do the math.

OpenStudy (anonymous):

never mind its d

OpenStudy (ipwnbunnies):

Not quite...

OpenStudy (ipwnbunnies):

(100cm/1m)^2 = 10000 cm^2 per m^2. 15*10000 = 150,000 cm^2